What is the mass of 1 mole of hydrogen, oxygen, carbon magnesium?

Grade:11

1 Answers

Leela Lochan
7 years ago
Mole=given mass /molecular mass. Molecular mass of the following elements is. Hydrogen -1. Oxygen -16.carbon-12.magnesium-24.using the above formulae we get mass =hydrogen mass is1×1. Oxygen=1×16.carbon=1×12.magnesium=1×24